College of Computer & Information Science Spring 2007 Northeastern University Handout 5 CSG250: Wireless Networks 13 February 2007

Problem Set 3 (due Friday, February 23)

(The problem numbers from the text are the same in both the first and second editions.)

1. (8 points) The DCF mode of IEEE 802.11

An ad hoc network using a 1 Mb/s IEEE 802.11 has 4 nodes: N1, N2, N3, N4. Assume that SIFS is 25 μs, PIFS 50 μs,, DIFS 100 μs, and slot time is 50 μs. Assume that at the beginning the channel is idle (no transmission), and that at instant 1, N2 has a packet to be sent to N4. At instant 2, both N1 and N3 have a packet to be sent to N4. Assume that the random number generator (for backoff) will give the following values for N1: 2, 5, …; for N2: 4, 3, …; and for N3: 1, 4, ….

Assume that we don’t use RTS/CTS or fragmentation, and that all data packets have the same length of 500 bytes and that the Ack packet has length 25 bytes. Furthermore the channel Bit Error Rate is assumed to be 0, and 802.11 provides the maximum possible throughput when there are no collisions. Show the execution of the DCF mode of IEEE 802.11.

2. (6 points) The PCF mode of IEEE 802.11

Suppose that four stations in an IEEE 802.11 infrastructure network are in a polling list. The stations transmit 20 ms voice frames produced by 64 kbps speec encoders. Suppose that the contention-free period is set to 20 ms. Sketch a point-coordination frame transfer with the appropriate values for interframe spacings, NAV, and data and ACK frames.

3. (6 points) Mobile IP

Problem 12.1.

4. (6 points) Efficiency of cellular networks

Problem 10.15.

5. (18 points) Access points and handoffs

Consider a network consisting of two access points: AP1 and AP2, located at positions (-100,

0) and (100, 0) on the plane, respectively, where the distance unit is meter. A mobile unit is communicating through this network, which offers a data rate of 1 Mbps, using BPSK modulation and operating at a frequency spectrum centered around 2.4 GHz. Each access point transmits at the level of 100mW and the noise density in the area is 10^{−16} W/Hz. Assume that all packets have the same size: 100 bytes.

(a) Draw a two-dimensional graph that (approximately) indicates the region in which the mobile incurs a frame error rate of at most 5% on transmissions from access point AP1.

(b) Suppose the mobile decides to switch from one access point to another whenever the signal from the newer access point is at least 20% stronger than from the first one. Draw a twodimensional graph that indicates the points at which a mobile will switch from AP1 to AP2.

(c) Repeat part (b) for the case when the mobile decides to switch whenever the signal from the newer access point is at least 20% stronger than from the first one and the frame error rate for communication with previous access point exceeds 5%. An approximate curve will suffice.

You may adopt a free-space loss model for idealized isotropic antenna (ignoring gains) and may use the following formula for the Bit-Error-Rate (BER) of BPSK modulation:

BER = ∫_{√\overline{2E_{b}/N_{0}}}^{∞} \frac{1}{\sqrt{2π}} e^{− \frac{u^{2}}{2}} du.

(To calculate the above numerically, you can use the complement of the error function used in statistical analysis – e.g., the ERFC function in Excel.)